OpenStudy (anonymous):

If f'(x)=(x+2)^2*(x-4). what would the value of x be where f has a local max. or min. and how would I find the point of inflection?

OpenStudy (anonymous):

equate (x+2)^2*(x-4)=0 and solve

OpenStudy (anonymous):

don't forget to use the second derivative

OpenStudy (anonymous):

what would the second derivative be

OpenStudy (anonymous):

f'(x)= (x+2)^2*(x-4)=0 f'' (x)= product rule and chain rule cuz of this \[ (x+2)^2*(x-4)\]

OpenStudy (anonymous):

now u know how to solve it

OpenStudy (anonymous):

so f"(x)= 2(2x^2 - 2x) ?

OpenStudy (anonymous):

@Goaliess1 almost correct 2(x+2)( 1)(x-4)+ (1)(x+2)^2

OpenStudy (anonymous):

would it be 3x^2 -12

OpenStudy (anonymous):

find x

OpenStudy (anonymous):

the question is no over yet

OpenStudy (anonymous):

we have x=2

OpenStudy (anonymous):

so would the local max be at x=2 and there is no local min? also would the points of inflection be x=+2 and -2?

OpenStudy (anonymous):

this shows first derivative f'(x)=(x+2)^2*(x-4)

OpenStudy (anonymous):

yes but can't you find the point of inflection by the change in concavity?

OpenStudy (anonymous):

(-2, 4) critical pts

OpenStudy (anonymous):

f(-2)=2(x+2)( 1)(x-4)+ (1)(x+2)^2 f(4)= 2(x+2)( 1)(x-4)+ (1)(x+2)^2 this is the inflection point will decided whether this eq is local min or max

OpenStudy (anonymous):

sorry about that got little bit confuse, really need to use the second derivative to find the point of inflection

OpenStudy (anonymous):

what @matricked is right f'(x)=(x+2)^2*(x-4) x=-2 x=4

OpenStudy (anonymous):

about the f(x) used it to find the global max and min

OpenStudy (anonymous):

sorry about early

OpenStudy (anonymous):

just finish solving this f(-2)=2(x+2)( 1)(x-4)+ (1)(x+2)^2 f(4)= 2(x+2)( 1)(x-4)+ (1)(x+2)^2 and finding if is loca max or min then u are done have awonderful day im off now